### v2

# Random forest–based out-of-distribution detection for robust lung cancer auto-segmentation

Abstract
Accurate detection and segmentation of cancerous lesions from computed tomography (CT) scans is essential for automated treatment planning and cancer treatment response assessment. Transformer-based models with self-supervised pretraining have achieved strong performance on in-distribution (ID) data but often generalize poorly on out-of-distribution (OOD) inputs. We investigate this behavior for lung cancer segmentation using an encoder–decoder model. Our encoder is a Swin Transformer pretrained with masked image modeling (SimMIM) on 10,432 unlabeled 3D CT scans spanning cancerous and non-cancerous conditions, and the decoder was randomly initialized. This model was evaluated on an independent ID test set and four OOD scenarios: chest CT cohorts (pulmonary embolism and negative COVID-19) and abdomen CT cohorts (kidney cancers and non-cancerous pancreas). OOD detection was performed at the scan level using RF-Deep, a random forest classifier applied to contextual tumor-anchored feature representations. We evaluated 920 3D CTs (172,650 images) and observed that RF-Deep achieved FPR95s of 18.26% and 27.66% on the chest CT cohorts, and near-perfect detection (<< 0.1% FPR95) on abdomen-CT cohorts, consistently outperforming established OOD methods. These results demonstrate that our RF-Deep classifier provides a simple, lightweight and effective approach for enhancing the reliability of segmentation models in clinical deployment.
keywords:
Swin Transformer, auto-segmentation, lung cancer, out-of-distribution
1 Purpose
Among deep learning architectures, models that combine pretrained transformer backbones with convolutional decoders have demonstrated high accuracies in volumetric auto-segmentation of organs and tumors from 3D radiological scans [Willemink2022, jiang2022self, NguyenAAAI2023, YanWACV2023, Qayyum2023BHI, gu2024build]. A major challenge in deploying these models at scale is the potential accuracy degradation when applied to real-world scenarios that differ from those seen during training. For example, models trained to segment malignant lung cancers from chest CTs may also encounter scans that have benign nodules from lung cancer screening or scans containing unrelated diseases such as pulmonary embolisms, producing unanticipated and incorrect results outside their intended scope.
Traditional segmentation metrics such as the Dice similarity coefficient (DSC) and Hausdorff distance at 95th percentile (HD95) are designed to assess performance on in-distribution (ID) datasets, but are insufficient indicators of models’ robustness to out-of-distribution (OOD) data. OOD detection methods relying on model confidence scores, such as MaxSoftmax [hendrycks17baseline], often fail in scenarios where models produce confidently incorrect segmentations [yeung2023calibrating, hendrycks2019anomalyseg, zimmerer2022mood]. Alternative approaches using secondary models like VQ-GANs [graham2022transformer, pinaya2022unsupervised] require large training data, are computationally intensive, and lack interpretability, limiting their practical utility in high-throughput clinical workflows. Radiomics feature-based methods [vasiliuk2023limitations, konz2024radmetricmedicalimage] have shown to be effective at distinguishing scans of organs but have not generalized yet to tumors. To address these limitations, we propose RF-Deep, a lightweight scan-level OOD detection framework that leverages tumor-anchored contextual features from backbone encoders for reliable clinical deployment.
We evaluated our approach on far-OOD scans (diseases in different anatomic sites) and challenging near-OOD scans (different diseases in the same site with similar appearance). Our contributions are:
  1. i.
    a lightweight, model–agnostic OOD detection framework using random forest on deep feature representations extracted from the segmentation model’s encoder. While random forests for OOD detection are established, our contribution lies in the task-anchored application, a practical deployment concern under-explored in medical imaging,
  2. ii.
    demonstration on lung cancer segmentation including near-OOD (pulmonary embolism, COVID-19) and far-OOD (kidney cancers, healthy pancreas) scenarios, and
  3. iii.
    a systematic comparison against established OOD detection methods using five public datasets comprising 172,650 images from 920 patient scans.
2 Method
2.1 OOD detection task definition
Let $\mathcal{D}_{\text{in}}$ denote the distribution of lung cancer CT scans, closely matching the lung CT data used to fine-tune our segmentation model, and $\mathcal{D}_{\text{out}}$ represents the scans that differ in pathology and anatomy. Our objective is to determine whether a new scan $x$ belongs to $\mathcal{D}_{\text{in}}$ or $\mathcal{D}_{\text{out}}$, at the scan level. We compute voxel-level metric scores, average them to obtain $S(x)$, and apply a decision threshold $\tau$: scans with $S\leq\tau$ are flagged as ID, and $S\geq\tau$ as OOD. Since tumor segmentation is binary, aggregation is restricted to the predicted tumor regions, leveraging model-relevant spatial context while excluding irrelevant anatomy.
2.2 Segmentation model architecture
Our segmentation model is a hybrid transformer–convolution architecture that combines the global contextual modeling of a hierarchical Swin Transformer [liu2021swin] encoder with the local spatial precision of a U-Net–based [ronneberger2015u] convolutional decoder to accurately delineate anatomical boundaries of tumors. The encoder uses a depth configuration of 221222-2-12-2 across four stages with $4\times 4\times 4$ patch size, enabling multi-scale feature aggregation, while windowed self-attention reduces computational complexity yet preserves global context for large volumetric inputs (128 $\times$ 128 $\times$ 128 voxels). The decoder was randomly initialized, whereas the encoder was pretrained using SimMIM [xie2022simmim], a self-supervised masked image modeling approach in which 75% of 3D patches are masked and reconstructed. This strategy leverages large-scale unlabeled CT data to learn robust, anatomy-aware feature representations [hatamizadehSwinUNETR2022, tang2022self, jiang2022self, jiang2024self] and provides a strong foundation for downstream tumor segmentation and OOD detection.
2.3 OOD Detection with RF-Deep
Patient scans exhibit high variability and unlike natural images, may span the large anatomical boundaries, making a global OOD assessment inefficient. We therefore focus on the most semantically relevant areas, the tumor regions predicted by the segmentation model. Our approach (Fig. 1) can be outlined in four steps: (i) fine-tune a segmentation model with pretrained backbone encoders for lung tumor segmentation, (ii) extract tumor-anchored 3D crops and multi-scale encoder features, (iii) aggregate features into fixed-length representation maintaining global context focused on clinically relevant tumor regions and train a random forest classifier (RF-Deep) with ID and OOD examples, and (iv) classify new scans as ID or OOD using the trained RF-Deep. Critically, RF-Deep adds negligible computational cost as the feature extraction stage leverages the already-stored encoder representations from segmentation, and random forest inference is orders of magnitude faster than running secondary deep models.
Figure 1: High-level pipeline of the proposed out-of-distribution (OOD) detection framework, focusing on segmentation for lung cancer CT scans. After fine-tuning, the encoder is frozen, and features are extracted from the patch embedding (PE) layer and Swin blocks (SB) across all four stages to be used with a random forest classifier.
2.4 Implementation details
Datasets. We fine-tuned the model on NSCLC Radiomics [aerts2015data] (N=317), with LRAD [bakr2017data] (N=140) as the ID dataset. OOD evaluation used RSNA-Pulmonary Embolism [colak2021rsna] (N=120), negative COVID-19 MIDRC [ricord2021covid] (N=120), KiTS-23 [heller2023kits21] (N=120), and Pancreas-CT [roth2015deeporgan] (N=82), totaling 920 scans.Segmentation. All models were implemented using PyTorch [paszke2019pytorch] and MONAI [cardoso2022monai], fine-tuned using cross-entropy and Dice loss with a batch size of 16 across 4 NVIDIA GPUs. We used a learning rate of $2\times 10^{-4}$, with linear warm-up and cosine annealing over 1000 epochs. Augmentations included flips, rotations, affine, and intensity shifts. Inputs were normalized (HU [400,400][-400,400]), resampled (1mm3), and cropped to $128^{3}$. Sliding window with 50% overlap was used for inference. RF-Deep. Random forest (1000 trees, max depth 20, balanced weights) via the scikit-learn library. Features were extracted from 8 tumor-anchored crops per scan, and predictions were averaged at inference. Comparative approaches. RF-Radiomics used 293 IBSI-compliant features [zwanenburg2020image] via PyCERR [deasy2003cerr, apte2018extension], reduced with recursive elimination, and with same config as RF-Deep. We also evaluated MaxSoftmax [hendrycks17baseline], MaxLogits [hendrycks2022logits], energy [liu2020energy], and entropy. OOD experiment protocol. RF-based approaches used a fixed 40/60 patient-level train–test split, repeated over 100 matched seeds to assess stability. Other baselines used the full cohorts as they require no auxiliary data. Evaluation metrics. We report AUROC and FPR95 (%) to measure ID–OOD separability [davis2006relationship, liang2017enhancing].
3 Results
Table 1: Out-of-distribution (OOD) detection performance across different OOD detection strategies for our segmentation model. Results are reported as AUROC ($\uparrow$) and FPR95 ($\downarrow$), with best values per dataset highlighted in bold.
Method Pulmonary embolism COVID-19 Kidney cancer Pancreas
AUROC FPR95 (%) AUROC FPR95 (%) AUROC FPR95 (%) AUROC FPR95 (%)
MaxSoftmax 88.74 37.01 89.53 53.25 96.72 12.34 97.26 14.29
MaxLogits 90.88 34.42 91.30 37.01 97.51 7.140 98.12 6.490
Energy 90.97 35.06 91.05 37.01 97.50 7.140 98.12 6.490
Entropy 90.59 33.81 92.47 53.24 96.97 12.23 97.55 12.95
RF-Radiomics 88.26 40.13 90.98 36.36 95.94 19.87 95.90 15.06
RF-Deep (Ours) 95.16 18.26 92.88 27.66 99.81 0.110 99.89 0.720
Figure 2: Representative results with green checkmarks indicating correctly detected OOD scans and the red cross highlighting missed cases. The red contours denote segmentations from the model, illustrating that naive confidence measures may fail when tumor-like structures exist in non-cancer cohorts.
Figure 3: Supporting analyses of our approach: (a) t-SNE visualization shows ID/OOD separation in encoder representations (the convex hull denotes extent of ID), (b,c) SHAP analysis highlighting feature importance and interpretability in RSNA-Pulmonary Embolism and KiTS-23 kidney cancer datasets respectively, and (d) Stage-wise ablation shows individual RF-Deep performance from Swin Transformer encoder features.
Our results demonstrate that RF-Deep provides robust scan-level OOD detection for lung cancer segmentation, outperforming established approaches including MaxSoftmax and MaxLogits ( 1). Encoder features from the SimMIM-pretrained and fine-tuned model, paired with RF, consistently yielded higher AUROC and lower FPR95 across all scenarios, with large margin gains in the abdominal cohorts (FPR95 $\leq$ 0.1, second best is 6 points away). The substantial performance differences, particularly in far-OOD scenarios, reflect the advantage of task-relevant contextual features over confidence-based heuristics. Fig. 2 visually shows RF-Deep’s ability to detect scans from completely different anatomical sites and different disease cases, with some limitations. In pulmonary embolism cases, we observed that the segmentation model often highlights tumor-like structures rather than the emboli; nevertheless, our contextual features enable RF-Deep to correctly classify these scans as OOD in most cases, surpassing non-contextual approaches. Additionally, we computed an unsupervised clustering of the features [masarczyk2024tunnel] across all cohorts (Fig. 3(a)) and observed semi-clear, distinct clusters for in-distribution (ID) and OOD cohorts, making them separable with the random forest classifier. Fig. 3(b) and Fig. 3(c) show the SHAP analysis on the RF-Deep providing interpretability and identifying which encoder features are most influential in classifying a scan as OOD, allowing for future work into feature reduction. Finally, stage-wise ablation (Fig. 3(d)), wherein only features corresponding to a particular stage of the transformer were used for training RF-Deep, demonstrated that mid- and early-deep level Swin Transformer features are most effective, aligning with our SHAP findings.
4 Conclusion
We performed a comprehensive evaluation of different OOD detection methods applied to the clinical task of lung cancer segmentation. Our analysis showed that, even though segmentation models can predict tumors incorrectly at times, RF-Deep, built on encoder features from our fine-tuned model with a random forest, consistently outperforms established methods. The approach is lightweight, interpretable, and improves the reliability of segmentation in both near- and far-OOD clinical scenarios. Future work involves scaling to more disease sites, understanding multi-class behavior, and evaluating generalization across other backbone pretraining strategies.
Acknowledgements.
This research was partially supported by NCI R01CA258821 and Memorial Sloan Kettering (MSK) Cancer Center Support Grant/Core Grant NCI P30CA008748.